# A Debiased Estimator for the Mediation Functional in Ultra-High-Dimensional Setting in the Presence of Interaction Effects

### v2

@@ Appendix A Proofs
$\mathbb{P}\left(\sqrt{n_{c}}\|\xi_{3}-\mu\|_{\infty}\geq 2C_{3}^{2}t\varsigma^{4}V_{2}S_{2}\mid A_{1:n}\right)$ $\leq\sum_{j=1}^{p}\mathbb{P}\left(\sqrt{n_{c}}\left(\xi_{3}-\mu\right)_{j}\geqslant 2C_{3}^{2}t\varsigma^{4}V_{2}S_{2}\mid A_{1:n}\right)$
$=\sum_{j=1}^{p}\mathbb{P}\left(\sqrt{n_{c}}t\left(\xi_{3}-\mu\right)_{j}\geqslant 2C_{3}^{2}t^{2}\varsigma^{4}V_{2}S_{2}\mid A_{1:n}\right)$
$=\sum_{j=1}^{p}\mathbb{P}\left({\sf exp}{(\sqrt{n_{c}}t\left(\xi_{3}-\mu\right)_{j})}\geq{\sf exp}{(2C_{3}^{2}t^{2}\varsigma^{4}V_{2}S_{2})}\mid A_{1:n}\right)$
$\leq\sum_{j=1}^{p}{\sf exp}{(-2C_{3}^{2}t^{2}\varsigma^{4}V_{2}S_{2})}\mathbb{E}\left[{\sf exp}{(\sqrt{n_{c}}t\left(\xi_{3}-\mu\right)_{j})}\mid A_{1:n}\right]$
@@
$\mathbb{P}\left(\sqrt{n_{t}}\|\xi_{3}-\mu\|_{\infty}\geq 2C_{3}^{2}t\varsigma^{4}V_{2}S_{2}\mid A_{1:n}\right)$ $=\mathbb{P}\left(\|\xi_{3}-\mu\|_{\infty}\geq 2C_{3}\varsigma^{2}\sqrt{2V_{2}S_{2}}\sqrt{\frac{\log{(p)}}{n_{c}}}\mid A_{1:n}\right)$
$\leq p^{-1}\,.$ (A.19)
Defining $K_{7}=C_{3}\varsigma^{2}\sqrt{2V_{2}s_{2}}$ and taking expectation with respect to to $\sigma$-field $\mathcal{F}_{0,n}$ on the both side of Equation (A.2), we conclude that $\|\xi_{3}-\mu\|_{\infty}\leq K_{7}\sqrt{\log(p)/n_{c}}$ with high probability.
@@
$\|\xi_{4}\|_{\infty}$ $\leq\left\|\mu_{c}\right\|_{\infty}\left|\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(\mathbf{X}_{c,i}-\mu_{c})^{\top}\Sigma_{c}^{-1}\mu\right|$
$=\left\|\mu_{c}\right\|_{\infty}\left|\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}P_{i}^{\top}\Sigma_{c}^{-1/2}\mu\right|$
$:=\left\|\mu_{c}\right\|_{\infty}\left|\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}Z_{c,i}\right|\,,$
where $Z_{c,i}=P_{i}^{\top}\Sigma_{c}^{-1/2}\mu$. By Chebychev’s inequality, we obtain:
@@
Observe that $\operatorname{Var}(Z_{c}\mid A_{i}=0)=\mu^{\top}\Sigma_{c}^{-1}\mu$ which is upper bounded by $V_{2}$ by Lemma A.1. Therefore, choosing $t=\sqrt{\log{(p)}/n_{c}}$, we have:
@@
Moreover, $\|\mu_{c}\|_{\infty}\leq C_{c}$ by Assumption 5.6. Hence, we conclude that $\|\xi_{4}\|_{\infty}\leq K_{8}\sqrt{\log{(p)}/n_{c}}$ with probability going to 1. Finally, combining this bound along with the upper bounds in Equation (A.13) and (A.15) we conclude that $\|a-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leq K_{5}\sqrt{\log{(p)}/n_{c}}$ with probability going to 1, for some constant $K_{5}>0$.
@@
Under the Assumption 5.6, where $\tilde{a}^{\top}\Sigma_{t}^{-1}\tilde{a}=\mathcal{O}(1)$ and $\mu^{\top}\Sigma_{c}^{-1}\mu=\mathcal{O}(1)$, we have $a^{\top}\Sigma_{t}^{-1}a=\mathcal{O}_{p}(1)$ and $\bar{X}^{\top}\Sigma_{c}^{-1}\bar{X}=\mathcal{O}_{p}(1)$. Also, we have $\check{a}^{\top}\Sigma_{t}^{-1}\check{a}=\mathcal{O}_{p}(1)$ where $\check{a}=[\mu\quad\hat{\mathbf{B}}_{0}\mu]$, i.e., for some constant $V_{1}^{\prime}$, $\check{a}^{\top}\Sigma_{t}^{-1}\check{a}\leq V_{1}^{\prime}$ with high probability.
@@
To prove the result of $a^{\top}\Sigma_{t}^{-1}a=\mathcal{O}_{p}(1)$ and $\bar{X}^{\top}\Sigma_{c}^{-1}\bar{X}=\mathcal{O}_{p}(1)$, where $a=[\bar{X}\quad\hat{\mathbf{B}}_{0}\bar{X}]$ , we will first focus on the behavior of $\bar{X}$ and $\hat{\mathbf{B}}_{0}$.
@@
by selecting $e^{t^{2}/2}=p^{\tilde{c}}$, where $\varphi_{c}$ is a compatibility constant in Assumption 5.9. The last line follows the union bound, Theorem 6.1 and the generalization of Lemma 6.2 in Bühlmann and Van De Geer, (2011). (Here, we generalize Lemma 6.2 in Bühlmann and Van De Geer, (2011) to the scenario where $(\Sigma_{c})_{jj}\leqslant S_{2}$ rather than the original condition $(\Sigma_{c})_{jj}=1$. This relaxation remains valid by appropriately setting $S_{2}=1$.). Hence, by leveraging the sparsity Assumption 5.4 and Assumption 5.6, where $\frac{q}{p^{\tilde{c}}}\rightarrow 0$ as $q\rightarrow\infty$, or equivalently, there exists $\tilde{c}$ such that $q\leqslant p^{\tilde{c}}$, we establish the following bound:
@@
Then, by applying Hoeffding’s inequality, we can deduce that, with high probability:
$\|\bar{X}-\mathbb{E}[X]\|_{\infty}\leqslant C_{X}\sqrt{\log(p)/n}.$ (A.21)
Consequently, we conclude that $a^{\top}\Sigma_{t}^{-1}a=\mathcal{O}_{p}(1)$ and $\bar{X}^{\top}\Sigma_{c}^{-1}\bar{X}=\mathcal{O}_{p}(1)$.
Similarity, as $\left\|\left(\mathbf{B}_{0}-\mathbf{\hat{B}}_{0}\right)\right\|_{1,\infty}=\mathcal{O}_{p}\left(s\sqrt{\log(p)/n_{c}}\right)$ we can conclude that for some constant $V_{1}^{\prime}$, $\check{a}^{\top}\Sigma_{t}^{-1}\check{a}\leq V_{1}^{\prime}$ with high probability. $\square$
@@
For notational convenience, in the following proof we drop the superscript of $U_{i}^{(0)}$ and $\epsilon_{i}^{(1)}$. Recall from the Proposition 4.1:
$\frac{\sqrt{n}(\hat{\theta}_{0}-\theta_{0})}{\sigma_{n}}$ $=\frac{\sqrt{n}}{\sigma_{n}}\left\{\left\|a-\mathbf{W}_{t}^{\top}\tau_{1}\right\|_{\infty}\left\|\hat{\phi}_{1}-\phi_{1}\right\|_{1}+\left\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right\|_{\infty}\|\hat{b}-b\|_{1}+\|\mathbf{B}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}\right\}$
@@
$\|\hat{\phi}_{1}-\phi_{1}\|_{1}=\mathcal{O}_{p}\left(k\sqrt{\frac{\log(p+q)}{n_{t}}}\right),\quad\|\hat{b}-b\|_{1}=\mathcal{O}_{p}\left(sk\sqrt{\frac{\log(p)}{n_{c}}}\right),\quad\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}=\mathcal{O}_{p}\left(k\sqrt{\frac{\log(q)}{n_{t}}}\right).$ (A.23)
The above risk bounds rely on the sub-gaussianity of the covariates and error terms (Assumption 5.9) and restricted eigenvalue (RE) condition of the sample covariance matrix. To obtain the rate of convergence of $\|\hat{\phi}_{1}-\phi_{1}\|_{1}$ and $\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$ we need $(\mathbf{W}_{t}^{\top}\mathbf{W}_{t})/n_{t}$ to satisfy the RE condition, and for obtaining the rate of convergence of $\|\hat{b}-b\|_{1}$, we need $(\mathbf{X}_{c}^{\top}\mathbf{X}_{c})/n_{c}$ to satisfy the RE condition. As their expected value is $\Sigma_{t}+\mu_{t}\mu_{t}^{\top}\succeq\Sigma_{t}$ and $\Sigma_{c}+\mu_{c}\mu_{c}^{\top}\succeq\Sigma_{c}$ and they satisfy the RE condition (Assumption 5.9), the sample second moment matrices $(\mathbf{W}_{t}^{\top}\mathbf{W}_{t})/n_{t}$, $(\mathbf{X}_{c}^{\top}\mathbf{X}_{c})/n_{c}$ also satisfy RE condition with high probability by Theorem 6 of Rudelson and Zhou, (2012). Next, we can use Corollary 2 of Negahban et al., (2012) for a column-standardized design matrix. Theorem 2.1 in Rudelson and Vershynin, (2013) can be applied to verify that our design matrices $\mathbf{W}_{t}$ and $\mathbf{X}_{c}$ are column-standardized with high probability,
$n_{t}^{-1/2}\|\mathbf{W}_{t,j}\|_{2}$ $\leq n_{t}^{-1/2}\|\mathbf{W}_{t,j}-\mu_{t}\|_{2}+\underbrace{\|\mu_{t}\|_{\infty}}_{\mathcal{O}(1)}\leq\left(\frac{c_{1}}{4}\iota^{2}\right)S_{1}^{1/2},\forall j=\{1,\ldots,p+q\},$ (A.24)
$n_{c}^{-1/2}\|\mathbf{X}_{c,k}\|_{2}$ $\leq n_{c}^{-1/2}\|\mathbf{X}_{c,k}-\mu_{c}\|_{2}+\underbrace{\|\mu_{c}\|_{\infty}}_{\mathcal{O}(1)}\leq\left(\frac{c_{2}}{4}\varsigma^{2}\right)S_{2}^{1/2},\forall k=\{1,\ldots,p\}.$
where the $\mathcal{O}(1)$ follows the Assumption 5.6. Hence, pairing these facts with the sparsity and sub-Gaussian noise assumptions in Assumptions 5.4 and 5.6, along with the condition stated in (A.24), Corollary 2 of Negahban et al., (2012) implies that $\hat{\phi}_{1}$, $\hat{b}$, and $\hat{\gamma}_{1}$, obtained by applying Lasso with parameters $(\lambda_{1},\lambda_{2})$ as specified in Theorem 5.11, satisfy the $\ell_{1}$-risk bound in Equation (A.23) with high probability.
@@
$=$ $\|\bar{X}-\mu+\mu\|_{\infty}\left[\max_{1\leqslant j\leqslant q}\left\|\left(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right)_{j}\right\|_{1}\right]\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
$\leqslant$ $(\underbrace{\|\bar{X}-\mu\|_{\infty}}_{o_{p}(1)}+\underbrace{\|\mu\|_{\infty}}_{\leqslant C_{X}})\left[\max_{1\leqslant j\leqslant q}\left\|\left(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right)_{j}\right\|_{1}\right]\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
@@
$\sqrt{n}\left\|a-\mathbf{W}_{t}^{\top}\tau_{1}\right\|_{\infty}\left\|\hat{\phi}_{1}-\phi_{1}\right\|_{1}+\sqrt{n}\left\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right\|_{\infty}\|\hat{b}-b\|_{1}+\sqrt{n}\|\mathbf{B}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
@@
where we use $n\sim n_{t}\sim n_{c}$. This concludes that $|\Delta_{n}|=o_{p}(n^{-1/2})$.
@@
$\kappa^{2}$ $\leq(\mathbf{W}_{t}^{\top}\tau_{1})_{j}^{2}\leq\|\tau_{1}\|_{2}^{2}\sum_{\{i:A_{i}=1\}}\mathbf{W}_{t,ij}^{2}\implies n\|\tau_{1}\|_{2}^{2}\geq\frac{\kappa^{2}}{\frac{n_{t}}{n}\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}\mathbf{W}^{2}_{ij}}\,,$
$\kappa^{2}$ $\leq(\mathbf{X}_{c}^{\top}\tau_{2})_{j}^{2}\leq\|\tau_{2}\|_{2}^{2}\sum_{\{i:A_{i}=0\}}\mathbf{X}_{c,ij}^{2}\implies n\|\tau_{2}\|_{2}^{2}\geq\frac{\kappa^{2}}{\frac{n_{c}}{n}\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}\mathbf{X}^{2}_{ij}}\,.$
@@
$n\|\tau_{1}\|_{2}^{2}\geq\frac{\kappa^{2}}{C_{1}},\ \ \ \ n\|\tau_{2}\|_{2}^{2}\geq\frac{\kappa^{2}}{C_{2}}\,.$
@@
This concludes the proof of the asymptotic ineligibility of the bias term.

### v3

@@ Appendix A Proofs
$\mathbb{P}\left(\sqrt{n_{c}}\|\xi_{3}-\mu_{X}\|_{\infty}\geq 2C_{3}^{2}t\varsigma^{4}V_{2}S_{2}\mid A_{1:n}\right)$
$\leq\sum_{j=1}^{p}\mathbb{P}\left(\sqrt{n_{c}}\left(\xi_{3}-\mu_{X}\right)_{j}\geqslant 2C_{3}^{2}t\varsigma^{4}V_{2}S_{2}\mid A_{1:n}\right)$
$=\sum_{j=1}^{p}\mathbb{P}\left(\sqrt{n_{c}}t\left(\xi_{3}-\mu_{X}\right)_{j}\geqslant 2C_{3}^{2}t^{2}\varsigma^{4}V_{2}S_{2}\mid A_{1:n}\right)$
$=\sum_{j=1}^{p}\mathbb{P}\left({\sf exp}{(\sqrt{n_{c}}t\left(\xi_{3}-\mu_{X}\right)_{j})}\geq{\sf exp}{(2C_{3}^{2}t^{2}\varsigma^{4}V_{2}S_{2})}\mid A_{1:n}\right)$
$\leq\sum_{j=1}^{p}{\sf exp}{(-2C_{3}^{2}t^{2}\varsigma^{4}V_{2}S_{2})}\mathbb{E}\left[{\sf exp}{(\sqrt{n_{c}}t\left(\xi_{3}-\mu_{X}\right)_{j})}\mid A_{1:n}\right]$
@@
$\mathbb{P}\left(\sqrt{n_{t}}\|\xi_{3}-\mu_{X}\|_{\infty}\geq 2C_{3}^{2}t\varsigma^{4}V_{2}S_{2}\mid A_{1:n}\right)$
$=\mathbb{P}\left(\|\xi_{3}-\mu_{X}\|_{\infty}2C_{3}\varsigma^{2}\sqrt{2V_{2}S_{2}}\sqrt{\frac{\log{(p)}}{n_{c}}}\mid A_{1:n}\right)$
$\leq p^{-1}\,.$
Defining $K_{7}=C_{3}\varsigma^{2}\sqrt{2V_{2}S_{2}}$ and taking expectation with respect to to $\sigma$-field $\mathcal{F}_{0,n}$ on the both side of Equation (LABEL:eq:bound_17), we conclude that $\|\xi_{3}-\mu_{X}\|_{\infty}\leq K_{7}\sqrt{\log(p)/n_{c}}$ with high probability.
@@
$\|\xi_{4}\|_{\infty}$ $\leq\left\|\mu_{X,c}\right\|_{\infty}\left|\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(X_{c,i}-\mu_{X,c})^{\top}\Sigma_{X,c}^{-1}\mu_{X}\right|$
$=\left\|\mu_{X,c}\right\|_{\infty}\left|\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}P_{i}^{\top}\Sigma_{X,c}^{-1/2}\mu_{X}\right|$
$:=\left\|\mu_{X,c}\right\|_{\infty}\left|\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}Z_{c,i}\right|\,$,
where $Z_{c,i}=P_{i}^{\top}\Sigma_{X,c}^{-1/2}\mu_{X}$. By Chebychev’s inequality, we obtain:
@@
Observe that $\operatorname{Var}(Z_{c}\mid A_{i}=0)=\mu_{X}^{\top}\Sigma_{X,c}^{-1}\mu_{X}$ which is upper bounded by $V_{2}$ by Lemma A.1. Therefore, choosing $t=\sqrt{\log{(p)}/n_{c}}$, we have:
@@
Moreover, $\|\mu_{X,c}\|_{\infty}\leq C_{c}$ by Assumption 5.7. Hence, we conclude that $\|\xi_{4}\|_{\infty}\leq K_{8}\sqrt{\log{(p)}/n_{c}}$ with probability going to 1. Finally, combining this bound along with the upper bounds in Equation (A.12) and (A.14) we conclude that $\|a-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leq K_{2}\sqrt{\log{(p)}/n_{c}}$ with probability going to 1, for some constant $K_{2}>0$.
@@
Under the Assumption 5.8, where $a^{*\top}\Sigma_{W,t}^{-1}a^{*}=\mathcal{O}(1)$ and $\mu_{X}^{\top}\Sigma_{X,c}^{-1}\mu_{X}=\mathcal{O}(1)$, we have $a^{\top}\Sigma_{W,t}^{-1}a=\mathcal{O}_{p}(1)$ and $\bar{X}^{\top}\Sigma_{X,c}^{-1}\bar{X}=\mathcal{O}_{p}(1)$. Also, we have, for some constant, $V_{1}^{\prime}>0$, $\tilde{a}^{\top}\Sigma_{W,t}^{-1}\tilde{a}\leq V_{1}^{\prime}$ with high probability.
@@
To prove the result of $a^{\top}\Sigma_{W,t}^{-1}a=\mathcal{O}_{p}(1)$ and $\bar{X}^{\top}\Sigma_{X,c}^{-1}\bar{X}=\mathcal{O}_{p}(1)$, where $a=(\bar{X}^{\top}\quad(\hat{\mathbf{B}}_{0}\bar{X})^{\top})^{\top}$ , we will first focus on the behavior of $\bar{X}$ and $\hat{\mathbf{B}}_{0}$.
@@
by selecting $e^{t^{2}/2}=p^{\tilde{c}}$, where $\varphi_{c}$ is a compatibility constant in Assumption 5.11. The last line follows the union bound, Theorem 6.1 and relaxation of Lemma 6.2 in Bühlmann and Van De Geer, (2011). (Here, we relax Lemma 6.2 in Bühlmann and Van De Geer, (2011) to the scenario where $(\Sigma_{X,c})_{jj}\leqslant S_{2}$ rather than the original condition $(\Sigma_{X,c})_{jj}=1$). Hence, by leveraging the sparsity Assumption 5.4 and Assumption 5.8, where $\frac{q}{p^{\tilde{c}}}\rightarrow 0$ as $q\rightarrow\infty$, or equivalently, there exists $\tilde{c}$ such that $q\leqslant p^{\tilde{c}}$, we establish the following bound:
@@
Then, with an application of Hoeffding’s inequality we can deduce with high probability:
$\|\bar{X}-\mu_{X}\|_{\infty}\leqslant C_{X}\sqrt{\log(p)/n}.$ (A.19)
Consequently, we conclude that $a^{\top}\Sigma_{W,t}^{-1}a=\mathcal{O}_{p}(1)$ and $\bar{X}^{\top}\Sigma_{X,c}^{-1}\bar{X}=\mathcal{O}_{p}(1)$.
Recall the definition of $\tilde{a}=[\mu_{X}\quad\hat{\mathbf{B}}_{0}\mu_{X}]$ in Lemma 5.12. Similarity, as $\left\|\left(\mathbf{B}_{0}-\mathbf{\hat{B}}_{0}\right)\right\|_{1,\infty}=\mathcal{O}_{p}\left(s\sqrt{\log(p)/n_{c}}\right)$ we can conclude $\tilde{a}^{\top}\Sigma_{W,t}^{-1}\tilde{a}\leq V_{1}^{\prime}$ with high probability. $\square$
@@
Recall from the Proposition 4.1:
$\frac{\sqrt{n}(\hat{\theta}_{0}-\theta_{0})}{\sigma_{n}}$
$=\frac{\sqrt{n}}{\sigma_{n}}\left\{\left\|a-\mathbf{W}_{t}^{\top}\tau_{1}\right\|_{\infty}\left\|\hat{\phi}_{1}-\phi_{1}\right\|_{1}+\left\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right\|_{\infty}\|\hat{b}-b\|_{1}+\|\mathbf{B}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}\right\}$
@@
$\|\hat{\phi}_{1}-\phi_{1}\|_{1}$ $=\mathcal{O}_{p}\left(k\sqrt{\frac{\log(p+q)}{n_{t}}}\right),$ (A.21)
$\|\hat{b}-b\|_{1}$ $=\mathcal{O}_{p}\left(sk\sqrt{\frac{\log(p)}{n_{c}}}\right),$
$\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$ $=\mathcal{O}_{p}\left(k\sqrt{\frac{\log(q)}{n_{t}}}\right).$
The above risk bounds rely on the sub-Gaussianity of the covariates and error terms (Assumption 5.11) and restricted eigenvalue (RE) condition of the covariance matrix. To obtain the rate of convergence of $\|\hat{\phi}_{1}-\phi_{1}\|_{1}$ and $\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$ we need $(\mathbf{W}_{t}^{\top}\mathbf{W}_{t})/n_{t}$ to satisfy the RE condition, and for obtaining the rate of convergence of $\|\hat{b}-b\|_{1}$, we need $(\mathbf{X}_{c}^{\top}\mathbf{X}_{c})/n_{c}$ to satisfy the RE condition. As their expected value is $\Sigma_{W,t}+\mu_{W,t}\mu_{W,t}^{\top}\succeq\Sigma_{W,t}$ and $\Sigma_{X,c}+\mu_{X,c}\mu_{X,c}^{\top}\succeq\Sigma_{X,c}$ and they satisfy the RE condition (Assumption 5.11), the sample second moment matrices $(\mathbf{W}_{t}^{\top}\mathbf{W}_{t})/n_{t}$, $(\mathbf{X}_{c}^{\top}\mathbf{X}_{c})/n_{c}$ also satisfy RE condition with high probability by Theorem 6 of Rudelson and Zhou, (2012). Next, we can use Corollary 2 of Negahban et al., (2012) for a column-standardized design matrix. Theorem 2.1 in Rudelson and Vershynin, (2013) can be applied to verify that our design matrices $\mathbf{W}_{t}$ and $\mathbf{X}_{c}$ are column-standardized ($X$ is column-normalized if $\|X_{j}\|_{2}/\sqrt{n}\leq C$ where $X_{j}$ is $j$-th column of $X$ (Negahban et al., , 2012)) with high probability,
$n_{t}^{-1/2}\|W_{t,j}\|_{2}$ $\leq n_{t}^{-1/2}\|W_{t,j}-\mu_{W,t}\|_{2}+\underbrace{\|\mu_{W,t}\|_{\infty}}_{\mathcal{O}(1)}\leq\left(\frac{c_{1}}{4}\iota^{2}\right)S_{1}^{1/2},\forall j=\{1,\ldots,p+q\},$ (A.22)
$n_{c}^{-1/2}\|X_{c,k}\|_{2}$ $\leq n_{c}^{-1/2}\|X_{c,k}-\mu_{X,c}\|_{2}+\underbrace{\|\mu_{X,c}\|_{\infty}}_{\mathcal{O}(1)}\leq\left(\frac{c_{2}}{4}\varsigma^{2}\right)S_{2}^{1/2},\forall k=\{1,\ldots,p\}.$
where the $\mathcal{O}(1)$ follows the Assumption 5.7. Hence, pairing these facts with the sparsity and sub-Gaussian noise assumptions in Assumptions 5.4 and 5.7, along with the condition stated in (A.22), Corollary 2 of Negahban et al., (2012) implies that $\hat{\phi}_{1}$, $\hat{b}$, and $\hat{\gamma}_{1}$, obtained by applying Lasso with parameters $(\lambda_{1},\lambda_{2})$ as specified in Theorem 5.13, satisfy the $\ell_{1}$-risk bound in Equation (A.21) with high probability.
@@
$=$ $\|\bar{X}-\mu_{X}+\mu_{X}\|_{\infty}\left[\max_{1\leqslant j\leqslant q}\left\|\left(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right)_{j}\right\|_{1}\right]\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
$\leqslant$ $(\underbrace{\|\bar{X}-\mu_{X}\|_{\infty}}_{o_{p}(1)}+\underbrace{\|\mu_{X}\|_{\infty}}_{\leqslant C_{X}})\left[\max_{1\leqslant j\leqslant q}\left\|\left(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right)_{j}\right\|_{1}\right]\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
@@
$\sqrt{n}\left\|a-\mathbf{W}_{t}^{\top}\tau_{1}\right\|_{\infty}\left\|\hat{\phi}_{1}-\phi_{1}\right\|_{1}+\sqrt{n}\left\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right\|_{\infty}\|\hat{b}-b\|_{1}$
$\qquad\qquad\qquad\qquad+\sqrt{n}\|\mathbf{B}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
@@
where we use $n\sim n_{t}\sim n_{c}$. This concludes $|\Delta_{n}|=o_{p}(n^{-1/2})$.
@@
$\kappa^{2}$ $\leq(\mathbf{W}_{t}^{\top}\tau_{1})_{j}^{2}\leq\|\tau_{1}\|_{2}^{2}\sum_{\{i:A_{i}=1\}}{W}_{t,ij}^{2}\implies n\|\tau_{1}\|_{2}^{2}\geq\frac{\kappa^{2}}{\frac{n_{t}}{n}\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}{W}^{2}_{ij}}\,$,
$\kappa^{2}$ $\leq(\mathbf{X}_{c}^{\top}\tau_{2})_{j}^{2}\leq\|\tau_{2}\|_{2}^{2}\sum_{\{i:A_{i}=0\}}{X}_{c,ij}^{2}\implies n\|\tau_{2}\|_{2}^{2}\geq\frac{\kappa^{2}}{\frac{n_{c}}{n}\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}{X}^{2}_{ij}}\,.$
@@
$n\|\tau_{1}\|_{2}^{2}\geq\frac{\kappa^{2}}{C_{1}},\ \ \ \ n\|\tau_{2}\|_{2}^{2}\geq\frac{\kappa^{2}}{C_{2}}\,.$ (A.27)
@@
Thus, we obtain $\|\tau_{1}\|_{2}=\|\tau_{2}\|_{2}=\Omega_{p}(1/\sqrt{n})$. This concludes the proof of the asymptotic ineligibility of the bias term.